OpenStudy (anonymous):

Two moles of compound P were placed in a vessel. The vessel was heated and compound P was partially decomposed to produce Q and R. A dynamic equilibrium between chemicals P,Q and R was established. At equilibrium x moles of R were present and the total number of moles present was (2+x/2). What is the equation for this equilibrium reaction? A) P <-> 2Q+R B) 2P <-> 2Q+R C) 2P <-> Q+R D) 2P <-> Q+2R

OpenStudy (anonymous):

You have 2 moles of P so P = 2P. You it was broken to produce 2 other compounds Q and R. They are in equilibrium. The total is represented as (2+x/2) And you have only 2 compounds Q and R.

OpenStudy (anonymous):

Also X amount of R was present at equilibrium.

OpenStudy (anonymous):

well thanx for sending details... what is the correct option then???

OpenStudy (abmon98):

Try each equation from the option 1)P-->2Q+R Initial: 2 - - At equilibrium x moles were present look at the ratio so P at equilibrium will be 2-x and Q will be 2x Add the total number of moles 2-x+2x+x=2-x+3x=2+2x 2)2P--->2Q+R initial 2 - - at equilibrium x moles were present x, look at the ratio so it will be 2-2x and Q will be 2x add them up. 2-2x+2x+x=2-x 3) 2P-->Q+R Initial 2 - - At equilibrium x moles of R were present. Look at the ration so P at equilibrium would be 2-2x and Q will be x. Add them up 2-2x+2x=2 4)2P-->Q+2R Initial 2 - - At equilbrium x moles of R are present, so P will be 2-2x and Q will be x/2. 2R=x R=x/2 if 1 mole is present add them up 2-x+x+x/2=2 2+x/2